1. Calculate the optical reflectivity and transmittivity (defined as the fraction of the incident power that is reflected and transmitted) of TM polarized light respectively when it is incident at 45 degrees relative to the surface normal of a glass window from air. The refractive index for air and glass are 1.5 and 1 respectively.
Answer #1

Fr the ven medivw gncidenk liðit ig. glaci S+28.13 N2 les = 28-13 Ro= 0.027g TsavzmitiviłThe required value of Reflectivity and transmitivity is obtained as in the box. The Frasnels equation is being used. Thank you. Please like the answer.